OpenStudy (anonymous):

For the simple harmonic motion equation d = 5sin(π/4 t) , what is the period?

jimthompson5910 (jim_thompson5910):

hint: the period of y = A*sin(Bx - C) + D is T = 2pi/B

OpenStudy (anonymous):

Okay so this is what I had got... I had to make sure: 2*π*f*t = π*t/4 f = 1/8 T = 1/f = 8

OpenStudy (anonymous):

THANKS!

jimthompson5910 (jim_thompson5910):

T = 8 is correct, nice work

OpenStudy (anonymous):

Then what about this one? For the simple harmonic motion equation d = 8sin(4πt) , what is the period?

jimthompson5910 (jim_thompson5910):

in this case, B = 4pi so T = 2pi/B = ???

OpenStudy (anonymous):

Can't be the same! Why does it look like the other one?

jimthompson5910 (jim_thompson5910):

the period formula is the same, but the value of B is different this time

OpenStudy (anonymous):

1/2 right?

jimthompson5910 (jim_thompson5910):

good, the period is 1/2

OpenStudy (anonymous):

Thanks! :)

jimthompson5910 (jim_thompson5910):

you got the hang of this

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

yes sir-ee!